OpenStudy (anonymous):

Find the value of y for a given value of x, if y varies directly with x. If y = 61 when x = –305, what is y when x = –315?

OpenStudy (anonymous):

y=kx 61=k (-305) k= -1/5 So, since y=kx and k is -1/5.. y= -1/5 (-315) y=63